# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 5.1. The classical Milnor-Moore theorem
[…]
induced by the forgetful functor $\mathrm{Alg}(\mathsf{Ch}_{\mathrm{K}})_{/{\mathbb{1}}}\to\mathrm{Alg}_{{\mathrm{Lie}}_{\mathrm{K}}}(\mathsf{Ch}_{\mathrm{K}})$. By rectification the functor (18) is equivalent to the forgetful functor
@@
so that by adjointness the functor (17) identifies with the functor $\theta_{!}.$ The functor (17) fits into a commutative square:
[…]
@@
(19) ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}\rightleftarrows{\mathrm{Hopf}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}:{\mathrm{Prim}}_{\geq 0},$
where ${\mathrm{Prim}}_{\geq 0}$ takes the connective cover of the primitives. By Proposition 5.35 (2) and (3) adjunction (19) restricts to an adjunction
@@
6. A forgetful functor to simplicial restricted Lie algebras
Notation 6.1.
For any field ${\mathrm{K}}$ let ${\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}^{\geq 0}$ be the $\infty$-category of connective ${\mathrm{H}}({\mathrm{K}})$-module spectra. Let ${\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}}$ be the category of restricted Lie algebras over ${\mathrm{K}}$ and $\mathsf{s}{\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}}$ the category of simplicial restricted Lie algebras over ${\mathrm{K}}$.
By Example 3.5 the forgetful functor $\mathsf{s}{\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}}\to\mathsf{s}\mathrm{Vect}_{\mathrm{K}}$ from simplicial restricted Lie algebras over ${\mathrm{K}}$ to simplicial vector spaces over ${\mathrm{K}}$ is monadic. By Lemma 7.12 it preserves sifted colimits. By [2, Theorem 5.1] this guarantees that $\mathsf{s}{\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}}$ inherits a right induced model structure from $\mathsf{s}\mathrm{Vect}_{\mathrm{K}}.$
In the latter section we constructed a forgetful functor
$\mathrm{Alg}_{{\mathrm{Prim}}{\mathrm{T}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}^{\geq 0})\to\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}^{\geq 0})$
from restricted $L_{\infty}$-algebras to Lie algebras in ${\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}^{\geq 0}$. In this section we factor this forgetful functor as
$\mathrm{Alg}_{{\mathrm{Prim}}{\mathrm{T}}}({\mathrm{Mod}}^{\geq 0}_{{\mathrm{H}}({\mathrm{K}})})\to(\mathsf{s}{\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}})_{\infty}\to\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}^{\geq 0}),$
where $(\mathsf{s}{\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}})_{\infty}$ is the $\infty$-category underlying the right induced model structure on the category $\mathsf{s}{\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}}$ of simplicial restricted Lie algebras over ${\mathrm{K}}$ (Theorem 6.7).
Notation 6.2.
For every category with weak equivalences $({\mathcal{C}},{\mathrm{W}})$ let ${\mathcal{C}}_{\infty}$ be the underlying $\infty$-category, i.e. the $\infty$-categorical localization with respect to ${\mathrm{W}}.$
We mainly use the latter notation when $({\mathcal{C}},{\mathrm{W}})$ has the structure of a model category or $({\mathcal{C}},{\mathrm{W}})$ is the category of cofibrant objects ${\mathcal{M}}^{\mathrm{cof}}$ in a model category ${\mathcal{M}}$ endowed with weak equivalences between cofibrant objects.

### v2

@@ 4.1. The classical Milnor-Moore theorem
(4.8) $\mathrm{Alg}(\mathsf{Ch}_{\mathrm{K}})_{/{\mathbb{1}}}[\text{quasi-isos}]\to\mathrm{Alg}_{{\mathrm{Lie}}_{\mathrm{K}}}(\mathsf{Ch}_{\mathrm{K}})[\text{quasi-isos}]$
induced by the forgetful functor $\mathrm{Alg}(\mathsf{Ch}_{\mathrm{K}})_{/{\mathbb{1}}}\to\mathrm{Alg}_{{\mathrm{Lie}}_{\mathrm{K}}}(\mathsf{Ch}_{\mathrm{K}})$. By rectification the functor (4.8) is equivalent to the forgetful functor
@@
so that by adjointness the functor (4.7) identifies with the functor $\theta_{!}.$ The functor (4.7) fits into a commutative square:
${\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Vect}}_{\mathrm{K}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}({\mathrm{Vect}}_{\mathrm{K}})_{/{\mathbb{1}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}_{\mathrm{Lie}}(\mathsf{Ch}_{\mathrm{K}})[\text{quasi-isos}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Alg}(\mathsf{Ch}_{\mathrm{K}})_{/{\mathbb{1}}}[\text{quasi-isos}].}$
@@
(4.9) ${\mathcal{U}}:\mathrm{Alg}_{\mathrm{Lie}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}\rightleftarrows{\mathrm{Hopf}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})})_{\geq 0}:{\mathrm{Prim}}_{\geq 0},$
where ${\mathrm{Prim}}_{\geq 0}$ takes the connective cover of the primitives. By Proposition 4.38 (2) and (3) adjunction (4.9) restricts to an adjunction
@@
4.2. Higher chromatic heights
Next we consider an application to chromatic homotopy theory. We fix the following terminology:
Let $p$ be a prime and $n\geq 1.$ We consider spaces and spectra localized at $p$, i.e. local for the homology theory associated to the $p$-local sphere spectrum. Let ${\mathrm{Sp}}_{(p)}\subset{\mathrm{Sp}}$ the full reflexive subcategory of $p$-local spectra and ${\mathcal{S}}_{(p)}\subset{\mathcal{S}}_{*}$ the full reflexive subcategory of $p$-local pointed spaces.
By [30, Theorem 2.2.] there are respective reflexive full subcategories
${\mathrm{Sp}}_{T_{n}}\subset{\mathrm{Sp}}_{(p)},\ {\mathcal{S}}_{v_{n}}\subset{\mathcal{S}}_{(p)}$
of $T_{n}$-local spectra and $v_{n}$-periodic homotopy types. By [5], [37] there is a functor $\Phi:{\mathcal{S}}_{(p)}\to{\mathrm{Sp}}_{T_{n}}$, the Bousfield-Kuhn functor, such that
  1. (1)
    The localization functor ${\mathrm{Sp}}_{(p)}\to{\mathrm{Sp}}_{T_{n}}$ factors as $\Phi\circ\Omega^{\infty}:{\mathrm{Sp}}_{(p)}\to{\mathrm{Sp}}_{T_{n}}$.
  2. (2)
    The localization functor ${\mathcal{S}}_{(p)}\to{\mathcal{S}}_{v_{n}}$ universally inverts the $\Phi$-equivalences, which are known as $v_{n}$-periodic equivalences.
Heuts [31] proves that the Bousfield-Kuhn functor $\Phi:{\mathcal{S}}_{v_{n}}\to{\mathrm{Sp}}_{T_{n}}$ lifts to an equivalence ${\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})$. Composing with the equivalence $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})\simeq\mathrm{Alg}_{{\mathrm{Lie}}}({\mathrm{Sp}}_{T_{n}})$ that loops the underlying spectrum, one obtains an equivalence $\Phi[-1]:{\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{{\mathrm{Lie}}}({\mathrm{Sp}}_{T_{n}})$.
We deduce the following chromatic variant of the Milnor-Moore theorem:
Theorem 4.40.
Let $n\geq 1$ be a natural and $X$ a $v_{n}$-periodic homotopy type. The unit
$\Phi(X)[-1]\to{\mathrm{Prim}}{\mathcal{U}}(\Phi(X)[-1])$
identifies with the Goodwillie completion $\Phi\to\lim_{n\geq 0}P_{n}(\Phi)$ evaluated at the loops $\Omega(X)$.
Proof.
By [30, Theorem 2.6.] the Bousfield-Kuhn functor $\Phi:{\mathcal{S}}_{v_{n}}\to{\mathrm{Sp}}_{T_{n}}$ lifts to an equivalence $\bar{\Phi}:{\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})$. In particular, $\Phi$ preserves limits and small filtered colimits. By [39, Remark 6.1.1.32.] this implies that the induced functor
$\Phi_{*}:\mathrm{Fun}({\mathcal{S}}_{v_{n}},{\mathcal{S}}_{v_{n}})\to\mathrm{Fun}({\mathcal{S}}_{v_{n}},{\mathrm{Sp}}_{T_{n}})$
sends the Goodwillie completion $\mathrm{id}_{{\mathcal{S}}_{v_{n}}}\to\lim_{n\geq 0}P_{n}(\mathrm{id}_{{\mathcal{S}}_{v_{n}}})$ of functors ${\mathcal{S}}_{v_{n}}\to{\mathcal{S}}_{v_{n}}$ to the Goodwillie completion $\Phi\to\lim_{n\geq 0}P_{n}(\Phi)$ of functors ${\mathcal{S}}_{v_{n}}\to{\mathrm{Sp}}_{T_{n}}$.
By [43, Theorem 11.3.] the Goodwillie completion of the identity of $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}})$ is
$\mathrm{id}\to\lim_{n\geq 0}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}(-).$
See also [38, Theorem 2.20.]. By stability the localization $L:{\mathrm{Sp}}\to{\mathrm{Sp}}_{T_{n}}$ preserves finite limits. By [39, Remark 6.1.1.32.] this implies that the Goodwillie completion of the identity of $\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})$ is
$\mathrm{id}\to\lim_{n\geq 0}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}(-),$
where the composition product and spectral Lie $\infty$-operad are in ${\mathrm{Sp}}_{T_{n}}.$
Since $\bar{\Phi}$ is an equivalence, by [39, Remark 6.1.1.30.] the induced functor
$\bar{\Phi}^{*}:\mathrm{Fun}(\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}}),\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}}))\to\mathrm{Fun}({\mathcal{S}}_{v_{n}},\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}}))$
sends the Goodwillie completion
$\mathrm{id}\to\lim_{n\geq 0}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}(-)$
to the Goodwillie completion
$\bar{\Phi}\to\lim_{n\geq 0}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}(-)\circ\bar{\Phi}$
of functors ${\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})$.
Since the forgetful functor $\nu:\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})\to{\mathrm{Sp}}_{T_{n}}$ preserves limits and small filtered colimits, by [39, Remark 6.1.1.32.] it sends the Goodwillie completion
$\bar{\Phi}\to\lim_{n\geq 0}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}(-)\circ\bar{\Phi}$
of functors ${\mathcal{S}}_{v_{n}}\to\mathrm{Alg}_{{\mathrm{Lie}}(1)}({\mathrm{Sp}}_{T_{n}})$ to the Goodwillie completion
$\Phi\to\lim_{n\geq 0}\nu\circ\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}(-)\circ\bar{\Phi}$
of functors ${\mathcal{S}}_{v_{n}}\to{\mathrm{Sp}}_{T_{n}}$. Consequently, by uniqueness of Goodwillie completions the Goodwillie completion
$\Phi\to\lim_{n\geq 0}\nu\circ\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}(-)\circ\bar{\Phi}$
of functors ${\mathcal{S}}_{v_{n}}\to{\mathrm{Sp}}_{T_{n}}$ identifies with $\Phi\to\lim_{n\geq 0}P_{n}(\Phi)$.
Hence for every $X\in{\mathcal{S}}_{v_{n}}$ the canonical map $\Phi(\Omega(X))\to\lim_{n\geq 0}P_{n}(\Phi)(\Omega(X))$ in ${\mathrm{Sp}}_{T_{n}}$ identifies with the map
$\Phi(\Omega(X))\simeq\Omega(\Phi(X))\to\lim_{n\geq 0}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}\bar{\Phi}(\Omega(X))\simeq\lim_{n\geq 0}\tau_{n}({\mathrm{Lie}}(1))\circ_{{\mathrm{Lie}}(1)}\Omega(\bar{\Phi}(X)).$
By Theorem 4.17 the latter is the unit $\Omega(\Phi(X))\to{\mathrm{Prim}}{\mathcal{U}}(\Omega(\Phi(X))).$